# Neutron and $\bm{\gamma}$-ray Discrimination by a Pressurized Helium-4 Based Scintillation Detector

### v1

@@ 3.4 Gamma-ray Response Systematics by G4
Based on the experimental and the simulation results on the $\gamma$-ray response of the PHe detector, we can conclude that: 1) the $E_{dep}\sim 280$ keV peak arising due to the most probable value of energy deposit by the recoiling electrons within the confined geometry of the PHe detector. This is also seen as rising bump at the higher channel numbers ($\sim 240$) in the experiment (see Fig. 8), apart from the spectral feature at the lower channel numbers ( $\leq\,150$) which has been demonstrated as arising from accompanying low energy $\gamma$-rays of the so-called room background. Based on the spectral profile of the $E_{dep}$ spectra, we have done fitting of the peak after subtraction of an exponential background for the $E_{dep}$ spectra shown in the Fig. 12, from 662 keV to 2.225 MeV $\gamma$-rays. The average $E_{dep}$ value for the peak was found to be: $282\pm 10$ keV. Based on the above and the fact that the $x$-axis of the plots of the Fig. 8 scale as $E_{dep}$, a first order calibration with calibration constant of 1.175 keV/channel on the $x$-axis is applied. 2) Observed attenuation of the low energy $\gamma$-rays by the HDPE layers, which is mostly due to scattering of the $\gamma$-rays, qualitatively agrees with the simulation. However, no quantitative comparison could be attempted in our study due to lack of information about the discrete $\gamma$-rays from the sources. 3) Production of the 2.225 MeV $\gamma$-rays and correlation with the attenuation of the fast neutrons by the HDPE layers have been found to demonstrate reasonable agreement between our experiment and the G4 simulation.
@@ 4 Scintillation signal estimates for the PHe detector
Figure 17: S3 vs. S1 for NR and ER events due to incident neutron
@@ 5 Conclusion
The experimental studies based on 252Cf fast neutron source, 137Cs $\gamma$-reference source and 241Am source have successfully demonstrated that a fast and a slow gate-based integration of the detector output pulses by the QDC is effective in achieving discrimination. Attenuation of the neutron band by the HDPE layers placed between the source and the detector also confirms the discrimination method followed. The charge contents of the pulses over the long gate effectively scale as the energy deposited $E_{dep}$ by the recoiling 4He in case of the neutrons, or the recoiling electrons in case of the $\gamma$-rays. Based on the estimated fast neutron energy spectrum arising from the spontaneous fission of 252Cf[12], and the PHe detector efficiency data[14] for the fast neutrons, the measured $E_{dep}$ spectral profile has been found to match reasonably well with theoretically estimated profile, folded by the detector efficiency. The lower $E_{dep}$ part of the spectral profile, however, do not match due to the difference in the threshold for our measurement and the PHe detector efficiency data, which was based on the time-of-flight measurements. The measured spectral distribution was calibrated over a range in terms of the neutron energy, based on the above match of the profiles. The same calibration was applied for the measured neutron spectra pertaining to the 241Am source based measurements.
@@
Interaction of the neutrons and the $\gamma$-rays with the HDPE layers was investigated in detail through Geant4 based simulation. Production of the neutrons from the 241Am source was demonstrated to be largely due to the ($\alpha,n$) reactions on the constituent elements of the glass substrate. Corresponding neutron spectrum was used for the event generation. Neutron attenuation by the HDPE layers was successfully reproduced to qualitatively match with the experimental results. Interaction of the $\gamma$-rays emitted from the 241Am source and also produced through neutron absorption by the HDPE layers with the pressurized helium active medium, were studied to find the origin of the peak-like structures in the measured spectra. It is found that a significant number of relatively higher energy $\gamma$-rays ($E_{\gamma}-\gtrsim 650\,{\rm keV}$) produce recoiling electrons, which dump around 280 keV of energy, resulting in the peak-like structure. Role of the SS housing of the detector and its geometry were also investigated as part of the systematic studies in support of the above.
@@
Figure 18: Sensitivity limit at 90% CL with no background
@@ 6 Acknowledgements
We are grateful to Prof. Satyaki Bhattacharya of SINP, India, for his invaluable discussions on the statistical analysis for the sensitivity calculation. We would also like to acknowledge technical and operational support from Chandranath Marick of SINP, India. One of us (SS) would like to acknowledge financial support from the Department of Atomic Energy Raja Ramanna Fellowship (DAE-RRF) scheme to carry out this work.

### v2

@@ 3.4 Gamma-ray Response Systematics by G4
Based on the experimental and the simulation results on the $\gamma$-ray response of the PHe detector, we can conclude that: 1) the $E_{dep}^{ee}\sim 280$ keV peak arising due to the most probable value of energy loss by the recoiling electrons within the confined geometry of the PHe detector. This is also seen as rising bump at the higher channel numbers ($\sim 240$) in the experiment (see Fig. 7), apart from the spectral feature at the lower channel numbers ( $\leq\,150$) which has been demonstrated as arising from accompanying low energy $\gamma$-rays of the so-called room background. Based on the spectral profile of the $E_{dep}^{ee}$ spectra, we fitted the peak with logistic function after subtracting an exponential background for the $E_{dep}^{ee}$ spectra shown in the Fig. 11, from 662 keV to 2.225 MeV $\gamma$-rays. The average $E_{dep}^{ee}$ value for the peak was found to be: $282\pm 10$ keV. Based on the above and the fact that the $x$-axis of the plots of the Fig. 7 scale as $E_{dep}^{ee}$, a first order calibration with calibration constant of 1.175 keV/channel on the $x$-axis is applied. 2) Observed attenuation of the low energy $\gamma$-rays by the HDPE layers, which is mostly due to scattering of the $\gamma$-rays, qualitatively agrees with the simulation. However, no quantitative comparison could be attempted in our study due to lack of information about the discrete $\gamma$-rays from the sources. 3) Production of the 2.225 MeV $\gamma$-rays and correlation with the attenuation of the fast neutrons by the HDPE layers have been found to demonstrate reasonable agreement between our experiment and the G4 simulation.
3.5 Neutron spectral calibration based on the neutron energy spectral data of 252Cf
It is demonstrated through our G4 simulation that the $\Delta E$ values of the energy deposition by the neutrons inside the detector have a linear dependence of the neutron energy $E_{n}$ (see Sec. 3.2 and Fig. 9). Based on this observation, an alternative energy calibration of the detector was attempted using the prompt neutrons from a 252Cf source. It is expected that the energy spectrum of the neutrons originating from the spontaneous fission is Maxwellian. However, after detailed analysis by various groups on the properties of neutron sources, and several meetings under the banner of International Atomic Energy Agency (IAEA), held during 1980 to 1987, the expert group on 252Cf spontaneous fission-based fast neutron sources, had proposed a corrected Maxwellian spectrum, based on experimental data and theoretical estimates[17] over various energy segments from 0.2 MeV to 20 MeV. Accordingly, the corrected Maxwellian spectrum $F(E_{n})$ as function of the kinetic energy ($E_{n}$) of the neutrons for the prompt fission neutrons is given by
A comparison of the experimentally obtained 252Cf prompt neutron spectrum and the one theoretically obtained as above, are shown in the Fig. 14. They are found to be in good agreement over the energy range of  0.7 MeV to  3.2 MeV. It is important to mention two major points here. a) The upper limit of the available energy range is due to the gain settings of the built-in amplifiers associated with the 4He detector. b) Our comparison, shown in the Fig. 14, manifests that the measured spectrum is constrained by a lower energy cut-off around 0.75 MeV, whereas the estimated spectrum, folded by the efficiency data, has a lower energy cut-off of around 0.35 MeV which arises entirely from the available intrinsic efficiency data in Ref. [19].
There are uncertainties in this method of calibration, which is primarily caused by the sparse data on the intrinsic efficiency. The measured and the simulated efficiency results considerably differ over various energy ranges[19]. Besides, interpolation procedure followed leaves scope for introducing additional uncertainties. Because of these uncertainties, we have adopted a linear calibration without a cut-off. However, a least square fit of the calibration data up to the quadratic term (ie. $E_{n}(x)=a_{0}+a_{1}x+a_{2}x^{2}$, where $x$ is the channel number), results in the following best fit values: $a_{0}=0.00167\pm 0.00212\,{\rm MeV},a_{1}=0.0125\pm 3.93\times 10^{-5}\,{\rm MeV/ch},a_{2}=1.87\times 10^{-8}\pm 1.48\times 10^{-7}\,{\rm MeV/ch^{2}}$. Relatively smaller values of $a_{0}$ and $a_{2}$ justify our choice of linear calibration.
It may be noted that the detector is also capable of detecting the $\gamma$-rays due to its electron recoil response, though the intrinsic detection efficiency may be very small compared to that of the neutrons. Though considerable details about the shape of the energy spectral distribution of $\gamma$-rays from spontaneous fission of 252Cf is available[20], similar energy calibration procedure could not be attempted for the $\gamma$-ray spectra, because of the overlap of the neutron and the $\gamma$-bands (see Fig. 2) at lower energies.
@@ 4 Scintillation signal estimates for the PHe detector
Figure 17: S3 vs. S1 for NR and ER events due to the incident neutrons.
@@ 5 Conclusion
The experimental studies based on 252Cf fast neutron source, 137Cs $\gamma$-reference source and 241Am source have successfully demonstrated that a fast and a slow gate-based integration of the detector output pulses by the QDC is effective in achieving discrimination. Attenuation of the neutron band by the HDPE layers placed between the source and the detector also confirms the discrimination method followed. The charge contents of the pulses over the long gate effectively scale as the energy deposited $E_{dep}$ by the recoiling 4He in case of the neutrons, or the recoiling electrons in case of the $\gamma$-rays. Based on the estimated fast neutron energy spectrum arising from the spontaneous fission of 252Cf[17], and the PHe detector efficiency data[19] for the fast neutrons, the measured $E_{dep}$ spectral profile has been found to match reasonably well with theoretically estimated profile, folded by the detector efficiency. The lower energy part of the spectral profile, however, do not match due to the difference in the threshold for our measurement and the PHe detector efficiency data, which was based on the time-of-flight measurements[19]. The measured spectral distribution was calibrated over a range in terms of the neutron energy, based on the above match of the profiles. The same calibration was applied for the measured neutron spectra pertaining to the 241Am source based measurements.
@@
Interaction of the neutrons and the $\gamma$-rays with the HDPE layers was investigated in detail through Geant4 based simulation. Production of the neutrons from the 241Am source was demonstrated to be largely due to the ($\alpha,n$) reactions on the constituent elements of the glass substrate. Corresponding neutron spectrum was used for the event generation. Neutron attenuation by the HDPE layers was successfully reproduced to qualitatively match with the experimental results. Interaction of the $\gamma$-rays emitted from the 241Am source and also produced through neutron absorption by the HDPE layers with the pressurized helium active medium, were studied to find the origin of the peak-like structures in the measured spectra. It is found that a significant number of relatively higher energy $\gamma$-rays ($E_{\gamma}\gtrsim 650\,{\rm keV}$) produce recoiling electrons, which dump around 280 keV of energy, resulting in the peak-like structure. Role of the SS housing of the detector and its geometry were also investigated as part of the systematic studies in support of the above.
@@
Figure 18: Projected sensitivity limit at 90% CL with no background; XENONnT upper limit at 90% CL (2023) result for reference [24]
@@ 6 Acknowledgements
We are grateful to Prof. Satyaki Bhattacharya of SINP, India, for the invaluable discussions on the statistical analysis for the sensitivity calculation. We would also like to acknowledge technical and operational support from Chandranath Marick of SINP, India. One of us (SS) would like to acknowledge financial support from the Department of Atomic Energy Raja Ramanna Fellowship (DAE-RRF) scheme to carry out this work.